### v1

# How to Train a Backdoor-Robust Model on a Poisoned Dataset without Auxiliary Data?

Abstract
Backdoor attacks have attracted wide attention from academia and industry due to their great security threat to deep neural networks (DNN). Most of the existing methods propose to conduct backdoor attacks by poisoning the training dataset with different strategies, so it’s critical to identify the poisoned samples and then train a clean model on the unreliable dataset in the context of defending backdoor attacks. Although numerous backdoor countermeasure researches are proposed, their inherent weaknesses render them limited in practical scenarios, such as the requirement of enough clean samples, unstable defense performance under various attack conditions, poor defense performance against adaptive attacks, and so on.
Therefore, in this paper, we are committed to overcome the above limitations and propose a more practical backdoor defense method. Concretely, we first explore the inherent relationship between the potential perturbations and the backdoor trigger, and the theoretical analysis and experimental results demonstrate that the poisoned samples perform more robustness to perturbation than the clean ones. Then, based on our key explorations, we introduce AdvrBD, an Adversarial perturbation-based and robust Backdoor Defense framework, which can effectively identify the poisoned samples and train a clean model on the poisoned dataset. Constructively, our AdvrBD eliminates the requirement for any clean samples or knowledge about the poisoned dataset (e.g., poisoning ratio), which significantly improves the practicality in real-world scenarios. Finally, we conduct extensive experiments and evaluate the effectiveness and superiority of the proposed AdvrBD in defending against $8$ state-of-the-art attacks (including $3$ adaptive attacks) on $4$ datasets (Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100). The experimental results demonstrate that our AdvrBD can achieve the average ASR with nearly $0.00\%$ in defending against $8$ state-of-the-art backdoor attacks while maintaining high accuracy of the main task. Compared with the $7$ state-of-the-art backdoor defense methods, our AdvrBD also performs more prominently. Furthermore, the excellent performance on $3$ different model architectures and poisoning ratios also highlights the remarkable generalization capability of AdvrBD.
Index Terms:
backdoor attack, backdoor defense, adversarial perturbation, model security.
I Introduction
Training deep learning models relies on a large amount of training data [1, 2, 3]. However, the training data are usually collected from the web or purchased from an untrustworthy third party, which leads to a risk of being poisoned or unreliable [4, 5]. For example, a company that wants to train a facial recognition model is likely to obtain poisoned data from a third party who aims to hijack the facial recognition model for profit [6, 7]. Therefore, the unreliable training dataset may bring many serious security threats to the model training side. One of the typical threats is the backdoor attack, which is usually conducted by poisoning the training data [8, 9]. In the backdoor attack, the attacker manipulates a few training samples by adding a specific trigger (e.g., a pixel patch or blending) and modifies the labels of these samples as a target class. The model trained on such a manipulated dataset can learn the main task while maintaining a backdoor function that can support the attacker in hijacking the model. The backdoored model maintains high accuracy on the normal samples but predicts the target class whenever the trigger pattern is attached to a test sample, as depicted in Fig.1. Backdoor attacks bring a significant security risk for model training owners as they are easily conducted and allow adversaries to control the trained model stealthily [10, 11]. For example, a backdoored traffic sign recognition model may misclassify a stop sign with the trigger patch as a no-tooting sign, which may cause a serious traffic accident. A face recognition model may misclassify an illegal user with the trigger as a valid user, which may pose a tremendous threat to the property of the valid user.
Figure 1: Overview of backdoor attacks.
In recent years, backdoor attacks have attracted much attention from researchers due to their serious threat to model security. Numerous studies focus on how to identify the poisoned samples and train a clean model on the poisoned dataset [12, 13, 14, 15, 16]. However, the existing defense methods have some non-negligible limitations in practical applications. (1) Requirement for clean samples. Their defense effectiveness heavily relies on obtaining a substantial number of clean samples [14, 15, 16], which is impractical for defenders in certain specific real-world scenarios [17]. (2) Unstable defense performance. Their actual defense performance varies significantly depending on the choice of backdoor attack methods and poisoning ratios [18, 13], which is usually unknown to defenders in advance in practical applications. (3) Inadequate defense against adaptive attacks. They may exhibit poor resilience against adaptive attacks [13], e.g., transforming some poisoned samples.
Therefore, in this paper, we aim to overcome the shortcomings mentioned above and propose a more practical and robust backdoor defense approach. To achieve these goals, there are three challenges that we need to address:
  • How to identify the poisoned samples regardless of the poisoning ratio and without the auxiliary clean dataset?
  • How to defend against various backdoor attacks and have a stable performance against adaptive attacks?
  • How to achieve satisfactory defense performance while maintaining a high accuracy of the main task?
To address the above challenges, we propose an Adversarial perturbation-based and robustness Backdoor Defense (AdvrBD) framework. Concretely, we first explore the intrinsic relationship between perturbation and backdoor through rich theoretical analysis and experimental verification. Based on the above exploration results, we roughly isolate the poisoned samples and clean samples by adding adversarial perturbation to the training data and then observing the change in the output probability distribution. Note that we just select some clean samples rather than precisely distinguish the poisoned samples in this step because of the certain overlap between the poisoned and the clean ones. Then, we train an enhanced backdoored model by unlearning the selected clean samples and relearning the remaining poisoned dataset. The poisoned samples and the clean samples can be precisely identified based on the enhanced backdoored model. Finally, we can train a clean model on the identified clean samples. To improve the performance of the clean model, we further fine-tune the clean model by relabeling the poisoned samples and adding them to the clean dataset. The main contributions are threefold:
  • Firstly, we discover an important observation that the poisoned samples present more robust to the potential perturbations than the clean samples. The theoretical analysis and experimental results also demonstrate our findings.
  • Then, we propose a novel and robust backdoor defense (namely AdvrBD) framework without any auxiliary clean dataset, which can resist various popular backdoor attacks and multiple adaptive attacks.
  • Finally, extensive experimental results demonstrate that when faced with $8$ state-of-the-art backdoor attacks on $4$ datasets, our AdvrBD outperforms the $7$ state-of-the-art backdoor defense methods in terms of decreasing the attack success rate (average ASR nearly $0.00\%$) while only incurring negligible accuracy loss. Moreover, our AdvrBD has a satisfactory generalization on multiple model architectures and different poisoning ratios.
The remainder of the paper is as follows. In Section II, we introduce the related works. Then, we will show the preliminaries and the threat model in Section III and Section IV, respectively. In Section V, we present the design of the proposed backdoor defense strategy. The experimental results are shown and analyzed in Section VI. Finally, we summarize in Section VII.
II Related Works
In this section, we review the relevant works of backdoor attacks and introduce the development of backdoor defense methods.
II-A Backdoor Attack Methods
In recent years, many backdoor attack approaches have been proposed, including patch-based attacks, visibility of trigger attacks, label consistency, and so on. Gu et al.[9] proposed a typical backdoor attack method by adding a specific patch on the samples and modifying the corresponding label as the targeted label. This attack achieves high attack success and has a low-performance impact on the main task. Chen et al.[19] proposed a backdoor attack in which the trigger is blending background rather than a single pixel. This means that the trigger is hard for human beings to notice. Liu et al.[20] proposed a more stealthy backdoor attack that plants reflections as a trigger into the victim model. This attack can be resistant to many existing defense methods. Liu et al.[21] proposed a lightweight backdoor attack that just needs to inverse the neural network to generate a general trojan trigger and fine-tune some layers of the network to implant the trigger. However, all the above backdoor attacks require poisoning the label, which is easier to detect. Therefore, some researchers have also proposed clean-label backdoor attacks. For example, Shafahi et al.[22] proposed a targeted clean-label poisoning attack. This attack crafts poison images that collide with a target image in feature space, thus making it undistinguishable from a network. Because the attacker does not need to control the label, it is more stealthy to conduct a backdoor attack. Turner et al.[23] proposed a clean-label backdoor attack based on adversarial examples and GAN-generated data. The key feature of this attack is that the poisoned samples appear to be consistent with their label and thus seem clean even from human inspection. Chen et al.[7] proposed an Invisible Poisoning Attack (IPA), which is difficult to detect by existing defense methods. This attack not only employs highly stealthy poison training examples with the clean labels (perceptually similar to their clean samples), but also does not need to modify the labels. Li et al.[24] proposed two stealthy backdoor attacks in which the triggers are derived from the covert features. Compared with the existing backdoor attacks, the trigger patterns of this attack method are invisible to human eyes. Moreover, it is difficult to recover the backdoor trigger through the optimization algorithm. Zhu et al.[25] proposed a transferable clean-label poisoning attack in which the poison samples are fabricated to surround the targeted sample in feature space. Saha et al.[26] proposed a novel backdoor attack in which the poisoned samples are similar to the clean samples with the correct labels. [27] and [28] proposed defense-resistant backdoor attacks in outsourced cloud environment.
II-B Backdoor Defense Methods
Many defense methods are proposed to resist the existing backdoor attacks. Some approaches require that the defender must own some reserved clean datasets. For example, Qi et al.[14] proposed a backdoor sample detection method that directly enforces and magnifies distinctive characteristics of the post-attacked model to facilitate poison detection. Guo et al.[15] proposed a universal detection approach based on clustering and centroids analysis. The approach can detect the poisoned samples based on density-based clustering and the clean validation dataset. Zhu et al.[29] proposed a defense method by inserting a learnable neural polarizer layer, which is optimized based on a limited clean dataset. The layer can purify the poisoned sample by filtering trigger information while maintaining clean information. For a backdoored model, Chen et al.[16] proposed a generic scheme for defending against backdoor attacks. The insight of this scheme is to localize the neuron set related to the trigger with the auxiliary clean dataset and suppress the compromised neurons. Ma et al.[30] proposed an input-level detection method. The intuition of this method is that even though a poisoned sample and a clean sample are classified into the target label, their intermediate representations are also different. Based on this observation, the poisoned samples can be detected easily. Wei et al.[31] presented a backdoor mitigation method using a small clean dataset. This method employs unlearning shared adversarial examples to purify the backdoored model. Researchers also proposed some backdoor defense methods without auxiliary clean datasets. It is a more strict and practical setting. For instance, Li et al.[12] proposed a defense approach that aims to train the clean model on the poisoned data. The main intuition of this method is that the models learn backdoor samples much faster than learning with clean samples. The backdoor examples can be easily removed by filtering out the low-loss examples at an early stage. Because the poisoned samples are much more sensitive to transformations than the clean samples in a backdoored model, Chen et al.[13] distinguish the poisoned samples from clean samples based on the feature consistency towards transformations. Weng et al.[32] found a trade-off between adversarial and backdoor robustness. Then, Gao et al.[33] challenged the trade-off between adversarial and backdoor robustness and proposed a backdoor defense strategy based on adversarial training regardless of the trigger pattern. Li et al.[34] found that the existing backdoor attacks have non-transferability. That is, the trigger sample is not effective in another model that has not been injected with the same trigger. Based on this observation, the authors proposed an input sample detection method by comparing the input sample and the samples picked from its predicted class label based on a feature extractor. Huang et al.[35] proposed a backdoor defense via decoupling the training process, thereby breaking the connection between the trigger and target label. Mu et al.[36] observed that the adversarial examples have similar behaviors as the triggered samples. Then, a progressive backdoor erasing method is proposed to purify the poisoned model via employing untargeted adversarial attacks. Tang et al.[37] presented a robust backdoor detection approach that can effectively detect data contamination attacks. Feng et al.[38] proposed a backdoor detection method for pre-trained encoders, requiring neither classifier headers nor input labels. Chen et al.[39] presented a robust backdoor defense scheme for federated learning. This scheme can overcome many backdoor attacks, including amplified magnitude sparsification, adaptive clipping, and so on.
Although many backdoor defense methods have been proposed, they have some limitations (e.g., requiring an auxiliary clean dataset) used in practical scenarios. Different from the existing backdoor defense methods, we first explore the relationship between the perturbation and the backdoor. Then, based on the exploration results, we plan to propose a robust backdoor defense method that does not require an auxiliary clean dataset and has a stable performance against various backdoor attacks on different models and poisoning ratios.
III Preliminaries
In this section, we introduce the mainly related technologies in this paper.
III-A Backdoor Attack
A backdoor attack is to inject a trigger into a model by poisoning the training dataset. During the inference period, the backdoored model performs well on the original task but outputs specific attacker-chosen answers when the input contains a specific trigger. For more clarity, we formalize the most common backdoor attack method BadNets [9] as follows:
Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a neural network for an image classification task. Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ be a clean training set where $\boldsymbol{x_{i}}\in\mathcal{X}$ and $\boldsymbol{y_{i}}\in\mathcal{Y}$ are the training image and the corresponding label, respectively. To conduct a backdoor attack, the attacker will choose a backdoor trigger $\boldsymbol{b}=(\boldsymbol{m},\boldsymbol{t})$ that consists of a blending mask $\boldsymbol{m}$ and a trigger pattern $\boldsymbol{t}$. In general, the trigger pattern is usually small just to achieve a stealthier backdoor attack. During the training process, the attacker randomly selects some training samples and poisons them by adding a specific backdoor trigger. For one poisoned training sample:
$\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}.$ (1)
where $\times$ is pixel-wise multiplication and $\boldsymbol{x}^{\prime}$ is the poisoned training sample. After modifying the training sample, the corresponding label is fixed with a target label. A successful backdoored model will maintain high accuracy on the main task while outputting the target label when the trigger $\boldsymbol{t}$ appears. After BadNets [9], many different attack variations have been proposed to enhance effectiveness and stealthiness. For example, Blend [19] and PhysicalBA [40] have designed more complex patterns. WaNet [41] has proposed a stealthier input-specific-trigger attack. SIG [42] has proposed a clean-label attack that is stealthier as it does not change the labels. To evade the existing defense method, some adaptive attacks (e.g., TaCT [37], AdaptiveBlend [43] and AdaptivePatch [43]) have also been proposed to improve the attack performance.
Figure 2: Threat Model of AdvrBD.
IV Threat Model
In this paper, we consider a practical scenario where a model owner aims to train a deep learning model based on the training dataset provided by an untrusted third party. However, the third party may poison the training dataset to hijack the model in the future. We assume that the model owner and the untrusted third party play the role of the defender and the attacker in the threat model, respectively. The threat model is shown in Fig.2. The exact goal, knowledge, and capabilities of the defender and the attacker are described in the following.
Defender’s goal: The defender aims to identify the poisoned samples and train a clean model on the poisoned dataset.
Defender’s knowledge and capabilities: The defender defines the model architecture, learning algorithm, and hyperparameters and trains the model. It has no knowledge about the poisoning ratio and without any auxiliary clean dataset.
Attacker’s goal: The attacker aims to inject a backdoor into the trained model by poisoning the training dataset.
Attacker’s knowledge and capabilities: The attacker can manipulate the training dataset, including poisoning a fraction of the training dataset and modifying the labels of the poisoned samples. However, the attacker can not access the model architecture and parameters. Moreover, it can not interfere with the model training process.
Figure 3: Framework of AdvrBD.
V Design of AdvrBD
In this section, we provide an overview of the proposed method and the detailed design.
V-A Overview of AdvrBD
In this paper, we focus on how to identify the poisoned samples and train a clean model on the poisoned dataset. It is not trivial to satisfy this goal since even a tiny number of poisoned samples can accomplish the backdoor insertion process. To tackle this problem, we explore the inherent relationship between the perturbation and the backdoor attack theoretically and experimentally at first. Based on the exploration results, we propose a three-stage backdoor defense method dubbed AdvrBD and the overall framework is shown in Fig.3. There are three main phases in the proposed AdvrBD. Firstly, we partition a tiny number of clean samples by using the difference between clean samples and poisoned samples in the face of perturbations. An enhanced backdoored model is trained by unlearning the chosen clean samples and relearning the remaining dataset. Then, we precisely distinguish the poisoned samples and clean samples with a high true positive rate based on the enhanced backdoored model. The samples that are misclassified by the enhanced backdoored model are defined as clean samples. The rest are the poisoned samples. A clean model can be trained based on the identified clean samples. Finally, we relabel the poisoned samples and combine them with the clean samples to obtain a completely clean dataset, which can be used to fine-tune the clean model to further improve the model performance. The design of AdvrBD is presented in detail as follows.
V-B Detailed Design of AdvrBD
In this section, we present the detailed design of our backdoor defense approach.
V-B1 Exploring the Relationship between Perturbation and Backdoor
To identify the poisoned samples accurately, we need to train an enhanced backdoored model by unlearning some clean samples. It’s vital to select a few clean samples without any auxiliary dataset and figure out how poisoned samples influence the prediction of backdoored models, through which we can design corresponding metrics to separate the clean samples and poisoned ones. We begin our exploration by reviewing the backdoor attack process first.
Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ be a clean training set and $C:\mathcal{X}\rightarrow\mathcal{Y}$ denotes the functionality of the target neural network. For each image $\boldsymbol{x}_{i}$ in $\mathcal{D}$, we have $\boldsymbol{x}_{i}\in\mathcal{X}=[0,1]^{C\times W\times H}$, and $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$ is the corresponding label, where $K$ is the number of label classes. To launch an attack, backdoor adversaries first need to poison the selected clean samples $\mathcal{D}_{p}$ with covert transformation $T(\cdot)$. Then the poisoned samples are mixed with clean ones before training a backdoored model, the process of which can be formalized as: $\mathcal{D}_{t}=\mathcal{D}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}={(x_{i}^{\prime},y_{t})|x^{\prime}=T(x),(x_{i},y_{i})\in\mathcal{D}_{p}}$. Many attacks have been proposed to make $T(\cdot)$ more stealthy and undetectable or make less assumption towards adversaries’ capability like poison ratio $pr=|\mathcal{D}_{p}|/|\mathcal{D}|$. Nevertheless, these adversaries hold the same goal to hijack the neural networks by training a malicious network:
$\min_{\boldsymbol{\theta}}\sum_{i=1}^{N_{b}}\mathcal{L}\left(f\left(\boldsymbol{x}_{i};\boldsymbol{\theta}\right),y_{i}\right)+\sum_{j=1}^{N_{p}}\mathcal{L}\left(f\left(\boldsymbol{x}_{j}^{\prime};\boldsymbol{\theta}\right),y_{t}\right).$ (2)
where $N_{b}=|\mathcal{D}|$ and $N_{p}=|\mathcal{D}_{p}|$.
[…]
$\mathcal{L}_{kl}(f(x),f(x\times(1-\hat{m})+\hat{m}\times\delta))).$ (3)
where $\delta=\max(\min(\frac{\nabla_{\delta}\ell(f_{\theta}(x),y)}{\|\nabla_{\delta}\ell(f_{\theta}(x),y)\|_{2}},\epsilon),-\epsilon)$ and $\hat{m}$ is a randomly-generated mask of the perturbation. As shown in Fig.5, on both BadNets and Blend attacks, most poisoned samples exhibit good consistency against adversarially generated perturbations $\delta$. To further explain this intriguing phenomenon, we propose the corresponding proof based on the recent studies [47] to analyze the characteristics of poisoned samples (detailed proof can be found in Appendix VIII):
Figure 4: Loss landscape of clean sample and poisoned sample.
Theorem 1. For $\mathcal{D}$, $\mathcal{D}_{p}$ that satisfy $N_{b}=|\mathcal{D}|$ and $N_{p}=|\mathcal{D}_{p}|$, i.i.d. sampled from an uniform distribution and belonging to $K$ classes. Assume that deep neural network $f(\cdot;\theta)$ is a multivariate kernel regression (RBF kernel) with the same objective as attackers. For a given attacked sample $\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}$, we have: $\lim_{N_{p}\rightarrow N_{b}}C\left(\boldsymbol{x}^{\prime}+\delta\right)=y_{t}$.
(a) BadNets Attack
(b) Blend Attack
Figure 5: Statistical histogram of training samples.
The theorem above demonstrates that when the size of $\mathcal{D}_{p}$ is close to that of the clean ones, poisoned samples present robustness towards potential perturbations. To figure it out intuitively, we also visualize the loss landscape of the clean and poisoned samples (Blend Attack), respectively, as shown in Fig.4. Note that the loss of the clean sample exhibits steeper changes when the perturbation is performed in the direction of the gradient. This observation also further confirms the previous theoretical analysis.
Algorithm 1 Untrusted Dataset Partition
0:  Untrusted Dataset $\mathcal{D}_{u}$, Backdoored Model $\hat{f}$, Perturbation Radius $\epsilon$, Patch Size $r$, Partition Rate $p$
0:  Partitioned Dataset $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$
1:  $Q=[\ ]$ // Initialize empty list $Q$
2:  for $(x,y)$  in  $\mathcal{D}_{u}$ do
3:     // Generate a mask of the perturbation patch
4:     $\hat{m}=RandomPatchMask(r)$
5:     $\delta=\max(\min(\frac{\nabla_{\delta}\ell(f_{\theta}(x),y)}{\|\nabla_{\delta}\ell(f_{\theta}(x),y)\|_{2}},\epsilon),-\epsilon)$
6:     $\hat{x}=x\times(1-\hat{m})+\hat{m}\times\delta$
7:     // Calculate $KL$ divergence of each sample
8:     $Q.append(\mathcal{L}_{kl}(f(x),f(\hat{x})))$
9:  end for
10:  // Partition $\mathcal{D}_{u}$ according to $Q$ and $p$
11:  $\mathcal{D}_{u},\mathcal{D}_{c}=sort(\mathcal{D}_{u},Q,p)$
12:  return  $\mathcal{D}_{u},\mathcal{D}_{c}$
Figure 6: ASR and ACC of the model during backdoor enhance training on CIFAR10 with ResNet18.
V-B2 Dataset Partition
Based on the findings in the above section, we plan to partition the untrusted dataset roughly by using the characteristic that the poisoned samples perform greater robustness against potential perturbations than the clean samples. However, as shown in Fig.5, these poisoned samples are still mixed with some clean ones. Therefore, in this section, we just pick out a small number of clean samples iteratively during the training stage rather than precisely selecting all the poisoned samples. Note that we do not require any auxiliary clean dataset.
Algorithm 1 shows the complete process of our dataset partition operation. This algorithm delineates a meticulous process for partitioning a dataset into trusted and potentially compromised subsets by exploiting the variances in the behavior of a model when subjected to perturbed inputs. At its core, the algorithm iterates over each sample $(x,y)$ in the dataset $\mathcal{D}_{u}$, applying a strategically-sized perturbation patch characterized by a radius $r$ and bounded by a perturbation radius $\epsilon$, to the input sample. This is achieved through generating a random patch mask $\hat{m}$ and calculating an optimal adversarial perturbation $\delta$, which is applied to the input $x$ to yield a perturbed counterpart $\hat{x}$. The key metric for partitioning, the KL divergence, is computed between the model’s predictions on the original and perturbed inputs, with these divergence scores being aggregated into a list $Q$. Subsequently, the dataset is partitioned into two subsets, $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$, based on the divergence scores relative to a partition rate $p$, indexing subsets with assumed clean and those potentially compromised, respectively. This method essentially leverages the sensitivity of backdoored models to perturbations in identifying and isolating suspicious data samples, thus serving as a robust strategy in the preprocessing phase of data handling. With the above steps, we obtain $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$. Here, $\mathcal{D}_{c}$ denotes the selected small number of clean samples. $\mathcal{D}_{u}$ represents the rest of the training data that contains the poisoned samples.
V-B3 Backdoor Enhancement and Standard Training
Note that a small number of clean samples $\mathcal{D}_{c}$ means nothing to train a clean model. However, motivated by [12], a small number of clean samples is enough to unlearn the functionality of a model on the main task. That’s to say, we can obtain a model $\hat{f}_{\theta}$ with a high attack success rate but low accuracy by unlearning some clean samples. To achieve a higher ASR and lower ACC, we train a backdoor-enhanced model iteratively by unlearning a small number of clean samples and learning the rest poisoned dataset. Concretely, for each iteration, $p$ of samples with the lowest consistency will be labeled as clean samples $\mathcal{D}_{c}$ and unlearned:
$\theta={\arg\max\limits_{\theta}}_{(x,y)\in\mathcal{D}_{c}}l(\hat{f}_{\theta}(x),y).$ (4)
During this process, more clean samples will be picked out for unlearning, and the model will be trained on a dataset with an increasing poison ratio. As illustrated in Fig.6, ASR and ACC of the model continue to increase and decline, respectively. Algorithm 2 shows the iterative backdoor-enhanced training process.
Algorithm 2 Backdoor Enhancement and Standard Training
0:  Untrusted Dataset $\mathcal{D}_{u}$, Backdoor Training Epochs $E_{b}$, Standard Training Epochs $E_{s}$
0:  Clean Model $f_{\theta}$
1:  $\mathcal{\hat{D}}_{u}=\mathcal{D}_{u}$ // Initialize $\mathcal{\hat{D}}_{u}$
2:  for $e$  in  $range(E_{b})$ do
3:     $\theta={\arg\min\limits_{\theta}}_{(x,y)\in\mathcal{\hat{D}}_{u}}l(\hat{f}_{\theta}(x),y)$
4:     // Update $\mathcal{\hat{D}}_{u}$ with partition algorithm
5:     // $\gamma$ is a coefficient
6:     $p=1-(e+1)\times\gamma$
7:     $\mathcal{\hat{D}}_{u},\mathcal{D}_{c}=Partition(\mathcal{D}_{u},\hat{f}_{\theta},p)$
8:     // Unlearn the clean samples
9:     $\theta={\arg\max\limits_{\theta}}_{(x,y)\in\mathcal{D}_{c}}l(\hat{f}_{\theta}(x),y)$
10:  end for
11:  // Pick out the clean samples
12:  $\mathcal{D}_{c}=\{(x,y)\in\mathcal{D}_{u}:\hat{f}_{\theta}(x)\neq y\}$
13:  for $e$  in  $range(E_{s})$ do
14:     $\theta={\arg\min\limits_{\theta}}_{(x,y)\in\mathcal{D}_{c}}l(f_{\theta}(x),y)$
15:  end for
16:  return  $f_{\theta}$
Obviously, our method does not end with training a backdoor-enhanced model, which plays a key role in identifying the poisoned samples. It’s natural to figure out the misclassified samples on the backdoored model and mark them as clean samples while the rest are labeled as poisoned samples. Finally, with the selected clean samples, we can train a clean model.
V-B4 Relabel and Relearn
To obtain a completed clean dataset and further improve the performance of the clean model, AdvrBD relabels the poisoned samples and merges them with clean samples to obtain a clean and complete dataset:
$\mathcal{D}_{com}=\mathcal{D}_{clean}\cup\{(x_{i},\hat{y}_{i})|\hat{y}_{i}=f(x),(x_{i},y_{i})\in\mathcal{D}_{p}\}.$ (5)
where $\mathcal{D}_{clean}$ denotes the selected clean samples based on the backdoor-enhanced model and $\mathcal{D}_{com}$ represents the complete clean dataset. Then, the complete dataset can be used to fine-tune the clean model to further improve its performance. Note that our experiment mainly shows the results before relabeling and relearning to follow the setting in previous works [12, 13], and the performance of the model fine-tuned on the complete dataset is given separately.
VI Experiments
In this section, we comprehensively evaluate the performance of the proposed AdvrBD. Firstly, we show the experimental settings, including the experimental environment, datasets, networks, backdoor attacks, and backdoor defenses. Then, we analyze and summarize the experimental results. Finally, we also present the ablation studies.
VI-A Experimental Settings
Experimental Environment: All experiments were conducted on a server equipped with Intel i9-9900K, 3.60GHz processor, 32GB RAM, NVIDIA GeForce RTX 3090, and PyTorch.
Datatsets: We perform our experiments on four image datasets, including Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100. The introduction of the above four datasets is shown as follows.
  • Imagnette [48]: Imagnette is a small dataset extracted from the large dataset ImageNet (more than $14$ million images, $20,000$ categories). Imagenette’s training set contains $9,469$ images, and its test set contains $3,925$ images, all in JPEG format. The image resolution is not uniform, but the width and height are no less than $160$ pixels.
  • Tiny ImageNet [49]: Tiny ImageNet contains $100,000$ images of $200$ classes ($500$ for each class) downsized to $64\times 64$ colored images. Each class has $500$ training images, $50$ validation images, and $50$ test images.
  • CIFAR10[50]: CIFAR10 dataset is a database of tiny images, including $50,000$ training images and $10,000$ testing images with $10$ classes.
  • CIFAR100 [50]: CIFAR100 contains $100$ categories. Each category contains $500$ training images with $32\times 32$ and $100$ test images with $32\times 32$.
Networks: We perform our experiments on three networks, including ResNet18, ResNet34, and MobileNetV2. It should be noted that unless otherwise specified, the default experimental network is ResNet18.
Backdoor Attacks: We employ $8$ state-of-the-art backdoor attacks to evaluate the proposed AdvrBD in our experiments, including three dirty-label attacks: BadNet [9], Blend [19], and PhysicalBA [40], one clean-label attack: SIG [42], one input-specific-trigger attack: WaNet [41], and three adaptive attacks: TaCT [37], AdaptiveBlend [43] and AdaptivePatch [43]. The detailed configurations of these backdoor attacks are shown in Table I.
TABLE I: The detailed settings of the $8$ state-of-the-art backdoor attacks.
Attack Method Poison Ratio Trigger Static/Dynamic Dirty/Clean Adaptive Source Target Cover Rate
BadNet 0.05 3*3 white square Static Dirty No / 1 /
Blend 0.05 HelloKity , alpha=0.2 Static Dirty No / 1 /
SIG 0.05 Sinusoidal signal Static Clean No / 1 /
WaNet 0.05 Warping-based triggers Dynamic Dirty No / 1 0.05
PhysicalBA 0.05 Firefox Static Dirty No / 1 /
TaCT 0.05 3*3 white square Static Dirty Yes 0 1 0.01
AdaptiveBlend 0.05 HelloKity , alpha=0.2 Static Dirty Yes / 1 0.01
AdaptivePatch 0.05 Firefox*4 Static Dirty Yes / 1 0.01
Backdoor Defense: We compare AdvrBD with $7$ established backdoor defense baselines, including FineTuning [51], FinePruning [51], CutMix [52], CLP [18], DBR [13], SCAnFT [37], and ABL [12].
Our detailed configurations for baseline defense (CIFAR10) are listed as follows:
  • FineTuning: We use $2,000$ reserved clean samples to finetune the “full layer” of the model, and the learning rate in the repair phase is set to $0.001$ with $10$ epochs.
  • FinePruning: We use $2,000$ reserved clean samples to prune the $30\%$ of the channels of the second layer.
  • CutMix: Probability of CutMix is set to $1$ with $\beta=1,\gamma=1$. To repair the backdoored model, the learning rate in the repair phase is $0.01$ for $10$ epochs.
  • CLP: Similarly, CLP needs to prune abnormal channels of the model that are out of $3$ standard deviations and repair the model for $10$ epochs.
  • DBR: DBR tries to pick out with high sensitivity to “rotate” and “affine” transformation. We set clean ratio and poison ratio to $0.4$ and $0.05$, respectively, and unlearn the poisoned samples for $5$ epochs with $0.0001$ unlearning rate.
  • SCAnFT: SCAnFT is a poisoned sample detection algorithm and here we adopt it to pick out the triggered samples, and unlearn them with the unlearning rate set to $0.0001$.
  • ABL: ABL tries to pick out poisoned samples by identifying the samples with lower loss values, and the threshold $\gamma$ of loss values is set to $0.8$, and the unlearning rate is $0.0001$ for $5$ epochs.
Evaluation Metrics: We employ the attack success rate (ASR) and accuracy (ACC) to evaluate the performance of the proposed backdoor defense framework. ASR indicates the ratio of the triggered samples that are misclassified as the target label. ACC represents the accuracy of the model on the clean samples. The lower ASR and the higher ACC indicate the better performance of the backdoor defense mechanism.
Specifically, we suppose a defense is successful if the post-defense ASR is under $20\%$, and unsuccessful otherwise, as done in prior works [14, 53].
TABLE II: The comparison results (%) between our AdvrBD and $7$ state-of-the-art backdoor defenses methods against $5$ state-of-the-art backdoor attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL AdvrBD
Dataset Types ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
Imagenette BadNet 94.10 93.70 71.47 32.13 61.50 54.85 72.30 56.00 33.24 6.09 69.81 57.62 46.54 16.34 57.61 58.45 73.15 0.00
Blend 96.60 95.50 70.08 88.92 60.94 93.63 55.12 92.80 13.30 0.00 45.15 70.64 57.89 17.17 62.60 79.22 73.13 0.00
SIG 96.40 86.30 70.36 71.47 60.39 15.51 70.63 88.09 16.90 74.24 61.50 0.28 41.83 0.28 16.90 83.38 73.96 0.00
WaNet 94.30 45.90 69.81 11.08 59.56 41.00 68.98 49.86 29.36 48.48 64.82 7.20 52.35 12.47 36.84 78.67 70.64 0.00
PhysicalBA 89.30 100.00 58.73 100.00 54.02 99.72 56.79 59.78 23.27 0.00 63.43 6.37 77.03 10.77 63.37 2.77 60.11 0.00
Average 94.14 84.28 68.09 60.72 59.28 60.94 64.76 69.31 23.21 25.76 60.94 28.42 55.13 11.41 47.46 60.49 70.20 0.00
CIFAR10 BadNet 92.00 95.20 83.30 96.81 75.05 93.19 81.21 83.30 34.18 0.55 67.47 1.21 69.89 2.09 71.98 0.00 85.72 0.00
Blend 84.50 99.90 84.62 100.00 79.23 98.79 80.44 99.67 51.00 3.52 37.03 98.46 78.57 1.43 82.20 5.71 85.71 0.00
SIG 84.70 95.40 82.31 95.93 80.88 90.11 79.12 99.12 37.47 71.97 38.57 93.52 66.92 3.63 85.93 48.35 87.69 2.64
WaNet 82.80 51.30 83.41 70.33 79.78 61.65 76.70 31.10 12.75 78.02 26.70 6.15 56.70 60.77 83.74 88.57 83.63 0.00
PhysicalBA 91.60 100.00 86.04 100.00 80.11 49.78 85.38 79.23 27.47 0.00 30.55 5.71 79.01 0.33 87.14 0.99 82.97 0.00
Average 87.12 88.36 83.94 92.61 79.01 78.70 80.57 78.48 32.57 30.81 40.06 41.01 70.22 13.65 82.20 28.72 85.14 0.53
CIFAR100 BadNet 72.10 98.20 58.40 91.40 42.00 91.09 42.91 86.03 54.55 83.30 29.35 0.10 56.78 0.00 63.16 0.10 60.53 0.00
Blend 51.90 100.00 60.12 99.80 31.58 99.79 48.99 99.70 34.31 59.51 8.20 89.88 48.89 0.00 61.40 1.21 60.53 0.00
SIG 54.40 95.00 60.32 96.86 40.28 67.71 48.99 95.95 43.72 69.13 7.39 80.10 44.03 2.94 62.96 90.59 57.49 1.11
WaNet 53.60 48.60 59.82 47.98 30.67 19.63 44.64 62.25 48.89 81.98 5.77 15.89 41.70 2.13 63.46 3.85 58.20 0.00
PhysicalBA 71.10 100.00 70.95 100.00 59.51 7.69 60.22 86.03 45.14 96.05 51.32 77.43 59.21 0.00 61.53 0.00 58.20 0.00
Average 60.62 88.36 61.92 87.21 40.81 57.18 49.15 85.99 45.32 77.99 20.41 52.68 50.12 1.01 62.50 19.15 58.99 0.22
TinyImageNet BadNet 37.80 97.95 46.00 98.54 26.87 97.58 31.55 90.84 5.08 84.90 0.70 0.00 32.21 0.00 46.80 0.00 35.73 0.00
Blend 34.60 99.95 49.20 100.00 22.29 96.98 32.11 99.80 10.37 18.32 21.54 99.75 5.33 0.00 38.65 0.00 35.73 0.00
SIG 36.50 98.80 47.26 97.79 25.52 89.43 33.87 99.39 7.60 100.00 0.75 0.00 29.89 0.00 26.27 78.86 38.60 0.00
WaNet 36.90 97.30 46.00 96.43 21.24 98.14 30.10 98.74 9.96 35.63 0.35 32.10 26.07 0.00 37.54 99.50 40.56 0.00
PhysicalBA 48.15 100.00 52.64 100.00 33.27 33.57 42.12 93.51 12.33 20.63 1.06 98.64 29.54 0.00 40.92 100.00 38.70 0.00
Average 38.79 98.80 48.22 98.55 25.84 83.14 33.95 96.46 9.07 51.90 4.88 46.10 24.61 0.00 38.04 55.67 37.86 0.00
VI-B AdvrBD Performance
VI-B1 Comparing to the Existing Defenses
In this paper, we employ $8$ state-of-the-art backdoor attacks (including $5$ familiar backdoor attacks and $3$ adaptive backdoor attacks) to demonstrate the effectiveness of our AdvrBD. In this section, we take into account $5$ state-of-the-art backdoor attacks and compare the performance of AdvrBD with $7$ other backdoor defense techniques. Table II shows the performance of the proposed AdvrBD method on Imagenette, CIFAR10, CIFAR100, and TinyImageNet. Obviously, our AdvrBD achieves the best results in reducing ASR against most backdoor attacks while maintaining a satisfactory ACC across all $4$ datasets.
Concretely, we mainly have five observations from Table II. The detailed analysis is shown as follows:
(1) Our defense method can resist to $5$ state-of-the-art backdoor attacks with an excellent performance. The ASR remains almost at $0.00\%$ on the $5$ backdoor attacks and $4$ datasets. The highest ASR is just $2.64\%$. Moreover, compared with no defense, our method has a slight drop on the main task on three datasets (CIFAR10: $87.12\%$ vs. $85.14\%$, CIFAR100: $60.62\%$ vs. $58.99\%$, and TinyImageNet: $38.79\%$ vs. $37.86\%$). However, we find that most defense methods fail to maintain the trade-off between ACC and ASR on ImageNette. This may be attributed to the small scale of this dataset since it is sampled from a larger and more complex dataset, ImageNet, which means that the convergence of the models trained on ImageNette may be largely affected even with little data missing. With 5% data poisoned, the original performance on this dataset will also be affected.
(2) Apart from SCAnFT, the other $6$ state-of-the-art backdoor defense techniques perform poorly. The highest average ASR is $98.55\%$ with the FineTuning on TinyImageNet. Only for the ABL on CIFAR100, the average ASR is $19.15\%$, which is less than $20.00\%$. The average ASR of all the other defense techniques is more than $20.00\%$. For SCAnFT, it has a fine defense performance on four datasets. The average ASR is $11.41\%$, $13.65\%$, $1.01\%$, and $0.00\%$, respectively, when the SCAnFT performs on the four datasets. However, SCAnFT has an obvious influence on the main task (from $94.14\%$ to $55.13\%$, from $87.12\%$ to $70.22\%$, from $60.62\%$ to $50.12\%$, from $38.79\%$ to $24.61\%$). This indicates that SCAnFT has significant limitations when used in practical scenarios.
(3) The existing defense methods perform unsteadily for the same backdoor attack on the different datasets. For example, for DBR under the SIG, the ASR on four datasets is $0.28\%$, $93.52\%$, $80.10\%$, and $0.00\%$, respectively. For ABL under the Blend, the ASR on four datasets is $79.22\%$, $5.71\%$, $1.21\%$, and $0.00\%$, respectively. Since the complexity, scale, and class numbers of the different datasets vary, the convergence speed of attack is also different. This results in a high false positive rate for poisoned sample identification when the backdoor task has not been learned. In this case, some clean samples may be picked out. Unlearning these clean samples will significantly influence the model performance.
(4) All the existing defense methods have an unstable performance against different backdoor attacks. They perform satisfactorily on some backdoor attacks and may be ineffective against the other backdoor attacks. This greatly reduces the availability of these defense methods in real-world scenarios because it is impractical for the defender to know the backdoor methods in advance. For example, ABL performs well on the BadNet attack but has an unsatisfactory performance on the SIG attack. Note that we have observed the different performance of these defense methods in different papers because their hyperparameters are different. We argue that it’s not practical to set different hyperparameters for different attacks because defenders have no knowledge about the potential attacks. Therefore, we use the same hyperparameter configurations for other defenses as well as AdvrBD.
(5) Compared with the $7$ state-of-the-art backdoor defense techniques, our AdvrBD can achieve the lowest ASR while maintaining a satisfactory performance on the clean samples. Moreover, the proposed AdvrBD has a more stable performance on different datasets and various backdoor attacks.
TABLE III: The comparison results (%) between our AdvrBD and $7$ state-of-the-art backdoor defenses methods against $3$ state-of-the-art adaptive attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL AdvrBD
Dataset Types ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
Imagenette TaCT 94.90 24.30 72.02 3.60 61.50 4.43 69.00 21.88 57.62 9.14 66.76 9.97 59.28 0.55 50.69 27.15 67.31 0.00
AdaptivaBlend 95.80 70.10 69.81 11.36 63.16 29.36 70.36 2.21 22.99 21.88 42.11 36.01 47.37 12.47 31.02 55.12 69.25 0.83
AdaptivaPatch 93.10 94.10 70.91 0.00 52.35 49.31 65.93 43.49 51.52 0.00 61.77 45.98 58.17 0.28 38.78 0.00 73.41 0.00
Average 94.60 62.83 70.91 4.99 59.00 27.70 68.43 22.53 44.04 10.34 56.88 30.65 54.94 4.43 40.16 32.92 69.99 0.28
CIFAR10 TaCT 82.10 63.20 82.63 65.71 75.05 65.71 78.68 41.43 37.91 3.08 77.58 63.19 72.42 62.20 85.16 48.02 82.42 0.00
AdaptivaBlend 84.00 53.30 82.09 66.48 80.88 54.40 74.95 49.12 49.45 5.49 45.27 12.09 78.90 0.22 88.68 3.41 83.30 0.00
AdaptivaPatch 84.80 37.30 84.07 49.78 82.42 0.11 61.65 6.48 70.33 2.09 80.22 25.27 72.64 0.00 83.30 99.34 82.86 0.00
Average 83.63 51.27 82.93 60.66 79.45 40.07 71.76 32.34 52.56 3.55 67.69 33.52 74.65 20.81 85.71 50.26 82.86 0.00
CIFAR100 TaCT 55.40 22.80 58.70 25.00 42.00 17.71 53.24 17.41 18.72 32.29 41.40 29.15 44.13 0.61 63.26 0.91 61.84 0.00
AdaptivaBlend 55.00 52.90 58.70 55.26 40.89 25.20 42.81 22.47 37.55 6.17 2.02 97.87 57.29 52.83 62.15 36.64 63.16 0.00
AdaptivaPatch 33.40 90.70 57.29 88.46 26.52 39.78 42.61 22.57 30.16 1.11 8.00 1.52 57.19 0.00 58.00 0.20 59.82 0.00
Average 47.93 55.47 58.23 56.24 36.47 27.56 46.22 20.82 28.81 13.19 17.14 42.85 52.87 17.81 61.14 12.58 61.61 0.00
TinyImageNet TaCT 36.95 13.65 46.55 10.87 26.07 8.05 32.36 19.38 2.62 0.00 0.55 0.00 25.42 0.00 40.71 20.73 38.80 0.00
AdaptivaBlend 34.75 73.40 46.00 74.53 24.81 35.78 32.91 56.06 18.52 20.23 0.20 77.10 23.45 0.00 30.20 52.04 40.82 0.00
AdaptivaPatch 38.55 24.35 47.10 99.90 22.40 32.56 32.56 53.85 6.95 8.81 0.45 0.55 33.67 0.00 42.78 99.70 39.71 0.00
Average 36.75 37.13 46.55 61.77 24.43 25.46 32.61 43.10 9.36 9.68 0.40 25.88 27.51 0.00 37.90 57.49 39.78 0.00
VI-B2 The Resistance to Adaptive Attacks
To further reveal the potential risk of the AdvrBD, we also consider the adaptive adversaries that try to design special backdoor attacks to escape our AdvrBD method. The adaptive adversaries deliberately establish dependencies between the backdoor and normal functionality, which immensely increases the difficulty of backdoor sample detection. Therefore, we employ three adaptive attacks (TaCT, Adap-Blend, and Adap-Patch) to demonstrate the effectiveness of our defense method. The three adaptive attacks build the dependencies between the backdoor and normal functions by using different poisoning strategies. Table. III shows the defense results of our method and other $7$ state-of-the-art backdoor defense techniques against the $3$ adaptive attacks on $4$ different datasets. From Table III, we have the following three observations. First, our AdvrBD has the lowest ASR (nearly $0.00\%$) while maintaining a satisfactory ACC of the main task. Then, some existing defense methods (CLP and SCAnFT) have a good defense performance against the three adaptive attacks, but they have an obvious decrease in ACC. Specifically, for CLP on the CIFAR10, the ASR is $3.55\%$, and the ACC drops from $83.63\%$ to $52.56\%$. Finally, our method has a better generalization on different datasets. For different datasets, our method always maintains a fine performance, while the performance of the other backdoor defense techniques has an obvious fluctuation. In summary, compared with the $7$ state-of-the-art backdoor defense techniques, our method has a better performance and greater generalization against adaptive attacks.
VI-C Ablation Studies
Different from the existing backdoor defense methods, we also propose to improve the performance of our AdvrBD further by relabeling and relearning the backdoor samples. Hence, in this section, we first show the AdvrBD’s performance after employing relabeling and relearning mechanisms. Then, to demonstrate our AdvrBD’s generalization, we also experimentally explore the performance of our AdvrBD on different partition rates, poison ratios, and models. Finally, we also explore the impact of our AdvrBD on the clean and poisoned dataset.
VI-C1 Performance with Relabeling and Relearning
In our AdvrBD, after identifying the backdoor samples and clean samples, we relabel the backdoor samples and merge them with clean samples to obtain a clean and complete dataset, which is used to fine-tune the clean model then. Accordingly, we conduct experiments on the ResNet18 and CIFAR10 datasets to show the effect of the relabeling and relearning mechanism. Table IV shows the ACC and ASR of our AdvrBD after using relabeling and relearning. Compared with Table II and III, we can know that the ACC further improves, and the ASR also decreases a little. Specifically, after employing relabeling and relearning, the average ACC and ASR changes from $84.29\%$ to $86.65\%$ and from $0.33\%$ to $0.03\%$ against the $8$ backdoor attacks, respectively. We infer that the model learns better due to the increase of clean samples. More interestingly, we find that employing relabeling and relearning has a positive effect on clean-label backdoor attacks (e.g., SIG). The ASR drops from $2.64\%$ to $0.22\%$ after equipping the relabeling and relearning mechanism. We deduce that the efficacy of backdoor task learning is closely linked to the initial weights of the model. Therefore, a pretrained model can hardly be backdoored by poisoned samples, which also indicates model service providers use pretrained models as the base models and fine-tune them with untrusted data to avoid poisoning attacks.
TABLE IV: The performance (%) of ResNet18 with relabeling and complete CIFAR10 dataset.
Stratagies→ W/O Relabeling and Relearning W/ Relabeling and Relearning
Attack↓ ACC ASR ACC ASR
BadNet 85.72 0.00 86.81 0.00
Blend 85.71 0.00 87.03 0.00
SIG 87.69 2.64 86.59 0.22
WaNet 83.63 0.00 86.92 0.00
PhysicalBA 82.97 0.00 83.52 0.00
TaCT 82.42 0.00 86.92 0.00
AdaptivaBlend 83.30 0.00 89.12 0.00
AdaptivaPatch 82.86 0.00 86.26 0.00
Average 84.29 0.33 86.65 0.03
VI-C2 Effectiveness with Different Partition Rates
In our AdvrBD method, we have to select a certain percentage of samples based on the $\mathcal{L}_{kl}$ value for training a backdoor enhancement model. This means that the partition rate may be important for AdvrBD’s performance. Therefore, to explore the effect of partition rate on the performance, we conduct experiments on CIFAR10 and $8$ backdoor attacks when the partition rate is $0.05$, $0.10$, $0.15$, $0.20$, and $0.25$, respectively. The experimental results are shown in Fig.7, which show that the partition rates have little influence on the AdvrBD’s performance. Our method on all the backdoor attacks apart from SIG has a stable performance. The ASR is always nearly $0.00\%$ except for SIG, in which the ASR causes a little fluctuation with a very small margin around $0.00\%$. We infer that it’s a normal fluctuation because of the occasionality of each experiment. For ACC, there is also a little fluctuation around $85.00\%$ for all the backdoor attacks. We infer that it is also a normal fluctuation due to the occasionality of each experiment.
Figure 7: The ASR and ACC (%) of AdvrBD with different partition rates.
TABLE V: Impact of poisoning ratio on defense performance (%).
Poison Ratio→ 0 0.1 0.15 0.2
Method↓ ACC ASR ACC ASR ACC ASR ACC ASR
BadNet 86.04 0.66 85.82 0.00 85.38 0.00 82.20 0.00
Blend 82.00 0.00 83.30 0.00 82.20 0.00
WaNet 83.74 0.00 85.05 0.00 82.42 0.00
PhysicalBA 82.42 0.00 81.76 0.00 81.10 0.00
TaCT 78.35 0.00 78.90 0.00 81.10 0.00
AdaptiveBlend 87.90 0.00 87.69 0.00 78.35 0.00
AdaptivePatch 87.58 0.00 87.25 0.00 86.26 0.00
Average 86.04 0.66 83.50 0.00 83.87 0.00 81.98 0.00
VI-C3 Effectiveness with Different Poisoning Ratios
In real-world application scenarios, it is impractical for the defender to access the poisoning ratios of training data. Therefore, we also demonstrate that our AdvrBD is suitable for multiple poison ratios. Here, we experiment it on CIFAR10 against $7$ backdoor attacks, including BadNet, Blend, WaNet, PhysicalBA, TaCT, AdaptivaBlend, and AdaptivaPatch with poisoning ratios up to $0.1$, $0.15$, $0.2$. Note that SIG is not considered here because it is a clean-label attack that can just poison one-class samples. Moreover, the number of one-class samples is a maximum of $10.00\%$ in CIFAR10. That is, the poisoning ratios of SIG on CIFAR10 do not exceed $10.00\%$. Table V shows the experimental results. From Table V, we can know that with the variation of poisoning ratios, the ASR of our method is always $0.00\%$ against the $7$ backdoor attacks. In particular, we can find that the ASR is $0.66\%$ when the poisoning ratio is zero. We suppose that the reason is the misclassification of the model on the clean samples. In addition, the experimental results show that our method also maintains a fine ACC against different backdoor attacks even without employing relabeling and relearning methods.
VI-C4 Effectiveness with Different Models
To demonstrate the generalizability of our method, we conduct experiments on ResNet34 and MobileNetV2 with the poisoning ratio of $0.05$. Table VI shows the experimental results on CIFAR10. From Table VI, we can find that our AdvrBD works similarly well on ResNet34 and MobileNetV2. The ASR is always $0.00\%$ in the two model architectures. Moreover, the ACC also maintains a satisfactory performance (from $88.60\%$ to $88.70\%$ in ResNet34, from $84.33\%$ to $81.89\%$ in MobileNetV2). The experimental results indicate that our AdvrBD has good generalizability across different models.
TABLE VI: The performance (%) of ResNet34 and MobileNetV2 on CIFAR10.
Model→ ResNet34 MobileNetV2
Attack↓ ACC ASR ACC ASR
No Attack 88.60 0.00 84.33 0.00
BadNet 88.35 0.00 80.30 0.00
Blend 88.35 0.00 81.76 0.00
SIG 89.56 0.00 82.31 0.00
WaNet 88.35 0.00 81.20 0.00
PhysicalBA 87.69 0.00 80.21 0.00
TaCT 89.45 0.00 82.63 0.00
AdaptiveBlend 89.01 0.00 80.22 0.00
AdaptivePatch 88.90 0.00 84.07 0.00
Average 88.70 0.00 81.89 0.00
VI-C5 Impact on clean and poisoned dataset
In real-world applications, it is impractical for the defender to know whether a given training dataset contains some poisoned samples. Therefore, to demonstrate the impact of our AdvrBD on the clean and poisoned dataset, we conduct experiments on CIFAR10 with the poisoning ratio of 00 and $0.05$. The experimental results (Table VII) illustrate that for training on the poisoned dataset, our AdvrBD can significantly reduce the average ASR (from $74.45\%$ to $0.33\%$) with a slight loss of ACC (from $85.81\%$ to $84.29\%$). On the other hand, when trained on a clean dataset, our AdvrBD has a minor impact on ACC (from $87.86\%$ to $86.04\%$). Note that the ASR is $0.66\%$ when training on the clean dataset with our AdvrBD. We infer that this can be attributed to the misclassification of the model on the clean samples. The experimental results further demonstrate that AdvrBD is a practical solution for dealing with unknown datasets.
TABLE VII: Impact of AdvrBD on the clean and poisoned dataset.
Attack Method Poison w/o defense Poison w/ defense Clean w/o defense Clean w/ defense
ACC ASR ACC ASR ACC ASR ACC ASR
BadNet 92.00 95.20 85.72 0.00 87.86 0.00 86.04 0.66
Blended 84.50 99.90 85.71 0.00
SIG 84.70 95.40 87.69 2.64
WaNet 82.80 51.30 83.63 0.00
PhysicalBA 91.60 100.00 82.97 0.00
TaCT 82.10 63.20 82.42 0.00
AdaptivaBlend 84.00 53.30 83.30 0.00
AdaptivaPatch 84.80 37.30 82.86 0.00
Average 85.81 74.45 84.29 0.33 87.86 0.00 86.04 0.66
VII Conclusion
In this paper, we first explore the relationship between the backdooor and perturbation by our theoretical analysis and experimental verification. Based on the exploration results, we propose a novel backdoor defense method (AdvrBD) to identify poisoned samples and train a clean model on a poisoned dataset. The proposed AdvrBD partitions the poisoned samples and clean samples based on their robustness to the adversarial perturbation. There is no requirement for any auxiliary clean dataset or knowledge about the poisoned dataset (e.g., poisoning ratios) in the AdvrBD. Extensive experimental results show the superior performance of AdvrBD in defending against $8$ state-of-the-art backdoor attacks. Compared with the $7$ advanced backdoor defense methods, our AdvrBD has a lower ASR while maintaining a satisfactory ACC on the main task. Besides, the experimental results also demonstrate that the proposed AdvrBD has fine generalization capability in different poisoning ratios and various model architectures. The relevant experimental results indicate our method has prominent potential and vital practicality for real-world application scenarios as well.